A rock is dropped vertically down from rest from the top of a 100-m high building. At what time and with what speed will the rock reach 50.0 m below the top of the building? (Ignore air resistance) A) 3.18 s, 31.3 m/s B 1.50 s, 19.8 m/s C) 4.36 s, 24.5 m/s D) 3.18 s, 11.6 m/s E) 9.80 s, 59.1 m/s
I know that the answer is a but what are the steps ??
Answer #1

Using the equation of motion:

s = ut + 0.5*at^2

Here, u = 0 as it starts off from rest

a = -9. 8 m/s2

So, -50 = 0 - 0.5*9.8*t^2

So, t = 3.18 s <----- answer

Now, using the equation of motion:

v = u + at

So, v = 0 + 9.8*3.18

So, v = 31.2 m/s <------ answer